OpenStudy (anonymous):

If 40 percent of a 20 gallon solution is alcohol,how many QUARTS of water must be added to make a new solution which is 25 percent alcohol?

OpenStudy (australopithecus):

we have a 40% alcohol solution so 0.40*20 = 8 gallons of alcohol in the 20 gallon solution we know 8g of alcohol/20g of total solution = 40% solution of alcohol in solvent so we set up the equation 25% = (8gallons of alcohol/x gallons of solvent)*100 0.25x = 8 x = 8/0.25 x = 32gallons 32gal - 20gal = 12gal thus we need to add 12 gallons of water to dilute it just convert using google or wolframalpha (seriously why aren't we using the metric system crazy americans)

OpenStudy (anonymous):

bases ten is overrated

OpenStudy (australopithecus):

I would probably kill myself if a lab I worked in didnt use the metric system, seriously 1 ml of water = 1g of water so handy

OpenStudy (australopithecus):

well under the right conditions obviously :)

OpenStudy (anonymous):

no one wanted to hear 28.35 grams of prevention is worth 453.6 grams of cure or her eyes are a blue million 1.609 kilometres

OpenStudy (australopithecus):

I wanted to hear :(

OpenStudy (australopithecus):

but in all seriousness what the are you blathering on about